OpenStudy (anonymous):

Please help: let f be the function given by a) Find the domain of f. b) Find f '(x) c) Find the slope of the line normal to the graph of f at x = 5 f(x)=((x^4)- (16x^2))^(1/2)

OpenStudy (amistre64):

chain rule; other than that its pretty basic

OpenStudy (amistre64):

say u = that middle stuff; then \[D\sqrt{u} = \frac{u'}{2\sqrt{u}}\]

OpenStudy (anonymous):

the original equation is all under a radical sign.. do i still set it up how i posted it?

OpenStudy (amistre64):

yeah, ^1/2 is fine for radicals

OpenStudy (amistre64):

some even perfer it when doing derivatives, helps them keep track of the "power" rule

OpenStudy (anonymous):

why is the denominator 2 radical u?

OpenStudy (amistre64):

becasue I cleaned it up by saying: u = that stuff in the middle ....

OpenStudy (amistre64):

\[D[u^{1/2}]=\frac{1}{2}u'u^{-1/2}\]

OpenStudy (anonymous):

oh . now i got it!

OpenStudy (amistre64):

fill in the x parts and solve it for x=5

OpenStudy (amistre64):

Find the slope of the line [normal] to the graph they try to trick you here

OpenStudy (anonymous):

the slope is y=mx+b ?

OpenStudy (anonymous):

I have another question: that chain rule, i tought is was au^a-1

OpenStudy (amistre64):

it is

OpenStudy (amistre64):

the slope IS f'(x) the normal is the slope perp to that: or simply, -1/f'(x)

OpenStudy (anonymous):

ok for the chain rule D[u^1/2]= 1/2 u^(-1/2)

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

well, since we are using u as a function of x; u(x) we pop out a derivative as well

OpenStudy (anonymous):

now i think i might be lost

OpenStudy (anonymous):

wait do you mean derivative of everything under the radical

OpenStudy (amistre64):

does u represent a function of x?

OpenStudy (amistre64):

u = x^4 - 16x^2 right?

OpenStudy (anonymous):

yeah im pretty positive thats under the radical

OpenStudy (amistre64):

then since u is a function of x and not just a variable in and of itself; anything done to u pops out its derivative the derivative rules SHOULD be taught as functions in their own right so that you dont have to "Add to" what youve already been taught with extra memory :)

OpenStudy (anonymous):

ok that makes much sense

OpenStudy (amistre64):

\[\frac{d}{dx}[u(x)]^n=nu(x)^{n-1}*u'(x)\]

OpenStudy (amistre64):

the chain rule is always working within these other rules .... always youre just not taught that since x' = 1 when it pops out

OpenStudy (amistre64):

\[D[x^2]=2x*x'\] but since x'=1 it seems rather redundant to do it that way

OpenStudy (anonymous):

ok i totally understand. now my answer will have no variables right?

OpenStudy (anonymous):

that is if i plug 5 into x

OpenStudy (amistre64):

itll have an x in there to plug in the 5; but yes, the slope itself and the normal produced will have no variables for a given point